OpenStudy (anonymous):

Classify the following equation: x2 + 4y2 - 2x + 16y + 13 = 0 Answer Parabola Circle Line None of the above

OpenStudy (anonymous):

x^2 + 4y^2 - 2x + 16y + 13 = 0 x^2 + 4y^2 - 2x + 16y = -13 (x - 1)^2 + 4(y + 2)^2 = 4 (x - 1)^2/4 + (y + 2)^2/1 = 1 It is none of the above because this is the form for an ellipse and not a circle of a line.